# An Empirical Study of Training ID-Agnostic Multi-modal Sequential Recommenders

### v2

@@ F.4. Mask Ratio of Images
F.5. Visualization of IDSR and MMSR
In this study, we approach SR as a unique downstream application of multi-modal learning. Specifically, we employ solely multi-modal representations to ascertain whether they could match or even supersede traditional item-ID representations. We visualize both MMSR and IDSR from two perspectives of item and user. For this purpose, we selected the Bili dataset, randomly choosing 1000 items and 10 users, and then leveraged t-SNE (Van der Maaten and Hinton, 2008) tool to unveil their latent distributions.
F.5.1. Visualization of Items
In our study, we examine the visual representation of items using item-ID, as presented in Figure 8. The insights gathered from this analysis revealed two prominent characteristics. Firstly, we observed a notable clustering of popular items. These items, having a high frequency of interaction, tend to dominate and overshadow the less prevalent items in the dataset. Secondly, we noticed that ’cold’ items, which are items with few interactions, tend to be scattered and emerge as outliers. These observed tendencies are problematic as they can impede the recommendation system’s learning efficacy. Specifically, the recommendation model may over-emphasize popular items, creating a bias, while simultaneously struggling to learn or incorporate ’cold’ items due to their sparse presence. This disparity can lead to a reduced overall quality of recommendations, especially for users with niche tastes.
On the other hand, our study also highlights the potential advantages of multi-modal representations. Notably, these representations exhibit a balanced distribution of items, showing little sensitivity to popularity variations. The key to this resilience lies in the multi-modal representation’s ability to tap into the semantic richness of an item’s text and imagery. By leveraging these additional data dimensions, multi-modal representations effectively mitigate the ’overheating’ phenomenon associated with popular items and counteract the ’cold’ issue linked with less interacted items.
Figure 8. Visualization of items in IDSR and MMSR.
F.5.2. Visualization of Users
In our research, we further conduct a comparative analysis between the user representations in IDSR and MMSR. Specifically, we observed the distribution of these representations within a given user and between different users. The results were illuminating.
The MMSR method reveals a tendency for tighter clustering of diverse items associated with individual users. This characteristic suggests that MMSR, with its multi-modal data, is proficient at discerning and capturing the transition patterns in a user’s activity. Such an attribute is instrumental in understanding and learning the behavior and preferences of users with fewer interactions or ’cold’ items, thus enhancing the overall recommendation quality. Simultaneously, despite achieving compact clusters for items within individual users, MMSR maintains pronounced distances between different users. This aspect illustrates the model’s capacity to uphold the uniqueness of each user’s preferences and interests, preventing the homogenization of user tastes.
This duality of facilitating personalized recommendations while supporting ’cold’ item learning sets MMSR apart. It leverages multi-modal data, enhancing the richness of representation and ensuring a balance between personalization and generalization, an essential aspect of any successful recommendation system. The efficacy of this approach is visually demonstrated in Figure 9, where the clustering and inter-user distances can be clearly observed.
Figure 9. Visualization of users in IDSR and MMSR

### v5

@@ 8.4 Mask Ratio of Images
8.5 Visualization of IDSR and MMSR
In this study, we approach SR as a unique downstream application of multi-modal learning. Specifically, we employ solely multi-modal representations to ascertain whether they could match or even supersede traditional item-ID representations. We visualize both MMSR and IDSR from two perspectives of item and user. For this purpose, we selected the Bili dataset, randomly choosing 1000 items and 10 users, and then leveraged t-SNE [57] tool to unveil their latent distributions.
8.5.1 Visualization of Items
In our study, we analyzed item visualizations by item-ID, as presented in Figure 4, revealing two main issues: a clustering of popular items and the scattering of less interacted, ’cold’ items, leading to a biased recommendation system favoring popular items. However, using multi-modal representations that include text and images leads to a more even distribution of items, reducing bias and improving the inclusion of ’cold’ items, thus enhancing recommendation quality.
Figure 4: Visualization of items in IDSR and MMSR.
8.5.2 Visualization of Users
In our study, we compared user representations in IDSR and MMSR, focusing on their distribution within and across users. MMSR showed a closer clustering of items for each user, indicating its effectiveness in recognizing user behavior and preferences, especially for those with ’cold’ items. Furthermore, MMSR preserved distinct distances between users, highlighting its ability to maintain individual user uniqueness and prevent preference homogenization. This is visually supported by Figure 5, which shows clear clustering and inter-user distances.
Figure 5: Visualization of users in IDSR and MMSR
9 Conclusion